OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. explain each step in figuring out how to determine the time it would take Jimmy to complete this job on his own.

OpenStudy (anonymous):

So far I have \[1/(x-8) + 1/x=1/3\] But whenever I solve for it, it gives me two different solutions, one negative and one positive. Do I just take the positive one because you can't have negative time?

OpenStudy (anonymous):

Because harry can rake 8 hours faster so I took jimmies time as x and subtracted 8 from it?

OpenStudy (anonymous):

Yeah but it doesn't say eight times as fast, it says eight hours faster. Doesn't that mean +/-8?

OpenStudy (anonymous):

you also need to make sure x - 8 > 0

OpenStudy (anonymous):

How do I do that?^^^^^^

OpenStudy (anonymous):

Guys please I need quick help

OpenStudy (anonymous):

the solution is 2 and 12, but 2 - 8 < 0, so 2 can't be the solution. thus 12 is the solution

OpenStudy (anonymous):

Morning, I'm not entirely sure about it being 8x

OpenStudy (anonymous):

it's not 8x. x-8 is correct

OpenStudy (anonymous):

anyway just take 12 as the solution

OpenStudy (anonymous):

I need to show my work though :'( I have 3x/3x(x-8)+3(x-8)/3x(x-8)=x(x-8)/3x(x-8) is that right so far

OpenStudy (anonymous):

I got the LCD and multiplied aeverything to get there

OpenStudy (anonymous):

Then I got rid of the denominator and I have this: 3x+3(x-8)=x(x-8)

OpenStudy (anonymous):

Never mind I got it.